calculate the de Broglie wavelength of a free electron with an energy of 50 GeV
1
Expert's answer
2021-09-01T11:20:44-0400

The de Broglie wavelength of this electron is


"\lambda=\frac{hc}{\sqrt{2Em_0c^2}}=2.48·10^{-17}\text{ m}."